OpenStudy (anonymous):

Part 1: Solve the following equation using one of the methods listed below. Part 2: Using complete sentences, explain why you chose the method you used. Equation: 0 = x2 + 10x + 5 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (anonymous):

Completing the square should do it

OpenStudy (anonymous):

But how?

OpenStudy (anonymous):

that equation cannot be factored so your only option is to complete the square of the quadratic formula... do you know how to do any of these two?

OpenStudy (anonymous):

*or the quadratic formula

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

i suggest quadratic formula since it is easier...

OpenStudy (anonymous):

quadratic formula is \(\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \) to use the quadratic formula, the equation must be written in this form: \[\large 0= ax^2+bx+c\] since it's already written in this form can you tell me what a= b= c= ??

OpenStudy (anonymous):

a = 1, b = 10 and c = 5 @dpaInc

OpenStudy (anonymous):

yes.. so all you gotta do now is put those values in to the quadratic formula.... notice you will have 2 solutions because of the plus/minus.....:) i can verify your answer if you want. just tell me your answer...:)

OpenStudy (anonymous):

@dpaInc -10 +or- √80 / 2

OpenStudy (anonymous):

yes... this is what you mean right? \[\large x=\frac {-10\pm\sqrt{80}}{2} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... that is correct... but does your teacher want it simplified?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

good...:)

OpenStudy (anonymous):

you can use your calculator to get a decimal approximation for that value....